OpenStudy (anonymous):

What is the vertex for the graph of y = 3x2 - 12x + 9? Answer (-2, 21) (-2, 45) (2, -3) (2, -9)

OpenStudy (anonymous):

formula of parabola: y=ax^2+bx+c vertex: (x,y) now, x=-b/2a with the given expression, x=-(-12)/2(3)=2 substitute the value of x to the expression, y=3(2)^2-12(2)+9=-3 vertex: (2, -3) :)